OpenStudy (princessashley):

The product of (3 + 2i) and a complex number is (17 + 7i) im confused can some one explain what i do i dont know werither to divide them or multipy them

OpenStudy (princessashley):

@shamim

OpenStudy (shamim):

Do u know i=√(-1)

OpenStudy (shamim):

i^2=-1

OpenStudy (princessashley):

yes

OpenStudy (princessashley):

i know ths i^2 is just -1

OpenStudy (princessashley):

so i get 37+45i

OpenStudy (princessashley):

thank you hopefully its right this is my second time taking this quiz

OpenStudy (shamim):

U r most welcome!!!!

OpenStudy (shamim):

Anyway which grade u r studying?

OpenStudy (princessashley):

12 grade liberal arts 2 and algebra 2

zepdrix (zepdrix):

Let's call our unknown complex number z, So we're told that (3+2i) times this complex number z, is equal to 17+7i,\[\large\rm (3+2i)z=17+7i\]To solve for z, we'll have to apply division,\[\large\rm z=\frac{17+7i}{3+2i}\]From there we will apply steps to put it into standard a+bi form. I hope those steps make sense, yes?

OpenStudy (princessashley):

so i used 3+2i as a conjugates right

OpenStudy (princessashley):

z=17+7i/3=2i * (3-2i)/(3-2i) right

zepdrix (zepdrix):

Multiplying by conjugate of your denominator? Yes good. So notice that the `numerator` that you end up with is the work that you already did! So that's good, we can use that for our numerator.

zepdrix (zepdrix):

Woops woops woops :) Nevermind, Our numerator is (17+7i)(3-2i) not (17+7i)(3+2i) like we had before. My mistake hehe

OpenStudy (princessashley):

im confused now...

OpenStudy (shamim):

Sorry i was wrong to understand the meaning of ur math

zepdrix (zepdrix):

\[\large\rm z=\frac{17+7i}{3+2i}\left(\frac{3-2i}{3-2i}\right)\]So we'll have some multiplication on the top and the bottom, \[\large\rm z=\frac{(17+7i)(3-2i)}{(3+2i)(3-2i)}\]

zepdrix (zepdrix):

On the bottom, do you remember what happens when you multiply complex conjugates? :)

OpenStudy (princessashley):

i foil them right

OpenStudy (shamim):

I think u know a^2-b^2=(a+b)(a-b)

OpenStudy (princessashley):

so i get 37/10 +45/10 i right

OpenStudy (shamim):

So (3+2i)(3-2i)=3^2-(2i)^2=?

zepdrix (zepdrix):

You `could` foil them, yes :) But it's good to remember how conjugates work. With real numbers, when you multiply conjugates, you end up with `the difference of squares`. Example: \(\large\rm (5-3)(5+3)=5^2-3^2=25-9\) With complex numbers, when you multiply conjugates, you end up with `the sum of squares`. Example: \(\large\rm (5-3i)(5+3i)=5^2+3^2=25+9\)

OpenStudy (princessashley):

okay im just going to move on im too confused .. but thank you

OpenStudy (shamim):

I can make it clearer if u give me time

OpenStudy (princessashley):

its okay i can miss a question i just need 4 out of 5

OpenStudy (shamim):

Can u tell me (2i)^2=?

OpenStudy (princessashley):

-4?

OpenStudy (shamim):

3^2-(2i)^2=?

OpenStudy (shamim):

(17+7i)(3-2i)=?

OpenStudy (shamim):

U know 3^2-(2i)^2=13 Right?

OpenStudy (princessashley):

13 and 65+7i

OpenStudy (shamim):

(17+7i)(3-2i) =51+14+21i-34i =65-13i

OpenStudy (princessashley):

okay

OpenStudy (shamim):

(65-13i)/13=? Right?

OpenStudy (princessashley):

7/13 i

OpenStudy (shamim):

No

OpenStudy (shamim):

65/13-i